OpenStudy (anonymous):

which equation represents a line that isperpendicular to a line that has slope -5? a. -x+5y=-10 b. -5x+y=1/5 c. y=-1/5x+4 d. y-10=5(x+1)

OpenStudy (anonymous):

I think it may be "c"

OpenStudy (amistre64):

c is close, but somethings awry with it

OpenStudy (amistre64):

perp slopes have the property that they they are flipped and negated

OpenStudy (amistre64):

-5/1 flips to -1/5 is close; but after negating we get: 1/5

OpenStudy (anonymous):

so it would be "a"

OpenStudy (amistre64):

a gets me all warm and fuzzy, it has to be a :)

OpenStudy (anonymous):

can you explain a reason

OpenStudy (amistre64):

sure given the form: ax+by=c lets transform this into something we are used to seeing: y = mx + intercept ax+by=c -ax -ax ----------- by = -ax + c ; divide off the b y = (-a/b) x + (c/b) since slope is defined conventionally as the letter m in the y=mx .... jargon m = -a/b we want -a/b = 1/5